OpenStudy (anonymous):

Find the differential dy for y =cos(pix) and evaluate it x=1/3 and dx=-.02 and round your results to 3 decimal places.

OpenStudy (anonymous):

\[\frac{dy}{dx} = y'(x) \rightarrow dy=y'(x)dx\]

OpenStudy (anonymous):

\[ y = \cos(\pi x) \rightarrow y'(x) = \pi\cos(\pi x) \]

OpenStudy (anonymous):

\[ \rightarrow dy = \pi \cos(\pi x)dx \]

OpenStudy (anonymous):

-pi sin(pi*x)

OpenStudy (anonymous):

oops, yeah..

OpenStudy (anonymous):

\[y = \cos(\pi x) \rightarrow y'(x) = -\pi\sin(\pi x)\] \[\rightarrow dy = -\pi \sin(\pi x)dx\]

OpenStudy (anonymous):

then plug in the x and dx correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i get dy to be .054 is there anyway to check if its correct